# Imperial College London, UK

### v1

@@ Abstract
Knee-less bipedal robots like SLIDER have the advantage of ultra-lightweight legs and improved walking energy efficiency compared to traditional humanoid robots. In this paper, we firstly introduce an improved hardware design of the bipedal robot SLIDER with new line-feet and more optimized mass distribution which enables higher locomotion speeds. Secondly, we propose an extended Hybrid Zero Dynamics (eHZD) method, which can be applied to prismatic joint robots like SLIDER. The eHZD method is then used to generate a library of gaits with varying reference velocities in an offline way. Thirdly, a Guided Deep Reinforcement Learning (DRL) algorithm is proposed to use the pre-generated library to create walking control policies in real-time. This approach allows us to combine the advantages of both HZD (for generating stable gaits with a full-dynamics model) and DRL (for real-time adaptive gait generation). The experimental results show that this approach achieves 150% higher walking velocity than the previous MPC-based approach.
Additional Video:https://youtu.be/B2tBiGhNZDM
Keywords: Bipedal Robot, Legged Locomotion, Hybrid Zero Dynamics.
@@ 1 Introduction
Bipedal robots have promising applications in human-centered environments including assistive care and agile locomotion. The vast majority of bipedal robots, e.g., [4][5][6], are built with full revolute joints—a design approach that introduces singularity issues when the knees are fully extended during walking. Additionally, knee actuators contribute significant weight and increase the inertia of the legs. To address this challenge, knee-less prismatic-jointed robots such as SLIDER[1] have been optimized with lightweight leg designs, which in turn significantly reduce the mass and inertia during walking. Due to its inherent linear structure that functions as an impact cushion, [7] employed Bayesian optimization to demonstrate the energy savings of the prismatic joint across a range of frequencies.
However, similar studies have lacked system stability considerations over the entire dynamic system, i.e., the planar contact constraints of the entire supporting foot are needed to allow the robot to track reference trajectories during its swing phase[9]. A static, precisely matched reference trajectory does not imply that the trajectory can still be optimal in the dynamic system. In fact, the study of the long-horizon of the state of a robotic system shows that even if a reference trajectory is initially optimal under static conditions, its performance degrades over time due to the accumulation of dynamic effects, uncertainties, and disturbances[11]. For robot locomotion, it is insufficient to design a controller solely on aligning the current state with the system’s linearized asymptotic stability point, especially when many robotic systems do not exhibit such a stability point at all[10].
@@ 2.1 Mass Reduction in Base Link and Child Joints
The subsequent child joints, including the yaw and roll joints, were redesigned with optimized, rigid 3D-printed parts featuring reinforcement ribs for enhanced strength in a compact form factor. Additionally, to minimize the radial load on the motor output shaft, both the roll and yaw axes adopt a coaxial mounting design. By integrating an additional bearing at the load’s distal end, the moment load on the motor’s internal output shaft bearing is reduced, improving motor lifespan and structural robustness. The overall assembly is illustrated in Figure 2(b), with the name of each joint axis labeled.
@@ 2.2 Differential Capstan Drive for Foot and Prismatic Sliding Joints Actuation
To adapt SLIDER to the Capstan drive mechanism, a differential drive scheme was proposed, utilizing two identical actuators to control both the foot joint and the sliding joint simultaneously. An illustration of how this differential mechanism works is shown in the figure 3 below.
@@
As depicted, each leg assembly features a tightened steel cable on both sides symmetrically, which wraps around a pulley mounted on an actuator, with the endpoints bolted to the feet(shown in figure 4). Although only one foot is necessary, the cable must extend to the top and be anchored to an output lever to ensure proper functionality of the differential capstan drive throughout the entire sliding span. For simplicity, the lower foot design is mirrored at the top. Additionally, the linear motion constrain of the leg is handled by a LML9B metal linear rail, which is tough, smooth, and small enough for this application.
@@
With the reduced Degree of Freedom (DoF) in the foot joint and updated controller framework, only linear contact is required. Therefore the new foot design is much more simplified, which can be done using a 12mm aluminum extrusion. The new model weighs only 73.7g, which is just 80.9% of the previous version’s weight.. Given the significant improvements outlined above, the new SLIDER design sets a benchmark for future hardware iterations and serves as a potential testing platform for the new control software framework. A complete CAD rendering and an assembled photo are shown in Figure 1 in the Introduction.
@@ 3.1 Extended Hybrid Zeros Dynamics
Here, the limit cycle is a closed trajectory in phase portrait. HZD cannot be directly applied to robots with prismatic joints, as the existence of a limit cycle is not guaranteed in such systems. Therefore, we propose the extended HZD method as shown in Fig.5
@@
Inside this figure, $S$ is the state space where the impact happened, and $Z$ is the invariant set based on Zero Dynamics assumption. $\Delta(\mathcal{S}\cap\mathcal{Z})$ means the discrete transformation when impact occurs. The left figure expresses the HZD applied in robot model with a virtual knee, and the right figure is the robot model with prismatic joints. The methods of mapping from left to right can be expressed by the following Fig.6
@@
In which the $q_{T}=[q_{\alpha}\ q_{\beta}\ q_{l}]^{T}$ express the position of hip pitch, ankle pitch, and sliding joint. $q^{\prime}_{T}=[q_{\alpha}\ q_{\beta}\ q_{l}]^{T}$, $q^{\prime}_{T}=[q^{\prime}_{\alpha}\ q^{\prime}_{\beta}\ q^{\prime}_{\theta}]^{T}$ express the position of hip pitch, ankle pitch and knee joint in virtual robot model. after the prismatic joint position transform $q_{l}=\sqrt{l_{1}^{2}+l_{2}^{2}+2l_{1}l_{2}cos(\theta)}$, equation (1) can be applied to transform all state portrait from virtual knee model to the prismatic model.
Thus, once we have obtained the virtual limit cycle with the knee, we can obtain the approximation of the limit cycle in the state portrait of the ideal locomotion gait of the corresponding prismatic joint model. In the method of using HZD in [10], we need to set up the virtual constraints; here, we refer to the setting style of [11] to make sure the output matrix is full-rank.
$L(\theta)=\sum_{s\in S}\frac{1}{T}\,\Bigl{(}\sum_{t=1}^{T}\Bigl{\|}\tau_{s}\bigl{(}{q}_{t},{\dot{q}}_{t}\bigr{)}\Bigr{\|}\Bigr{)}.$ (2)
We have used only the equation (2) as the cost function to optimize the energy consumption. Here $T$ is the time period for each step, and $\tau$ is the torque output by the control law generated by the HZD method.
@@ 3.2 Guided DRL structure
Since in the simulation, the coupling of our sliding joint and ankle joint is difficult to achieve, as shown in Section 2, we directly treat the upper foot and the lower foot as independent actuated joints while making the torque output of the upper foot always 0. We use the proximal policy optimization(PPO) algorithm with eHZD gait library for control policy training, and the whole training framework of Guided DRL is shown in Fig.7
@@
Here, the number of environments is $n_{env}$. The observation space is $O_{t}\subseteq\mathbb{R}^{n_{env}\times 235}$, which consists of CoM linear velocity space $v_{lin}\subseteq\mathbb{R}^{n_{env}\times 6}$, CoM angular velocity space $v_{ang}\subseteq\mathbb{R}^{n_{env}\times 6}$, weight projection space $g_{pro}\subseteq\mathbb{R}^{n_{env}\times 3}$, CoM velocity command space $C\subseteq\mathbb{R}^{n_{env}\times 4}$, actuated state position space $q_{dof}\subseteq\mathbb{R}^{n_{env}\times 12}$, actuated state velocity space $\dot{q}_{dof}\subseteq\mathbb{R}^{n_{env}\times 12}$, action space $A_{t}\subseteq\mathbb{R}^{n_{env}\times 12}$, and a height point cloud $H_{t}\subseteq\mathbb{R}^{n_{env}\times 180}$. The actions space is the desired position space for all actuated joints, which will be mapped to the final motor torque space by a PD controller. Here, the Actor network is composed of 3 hidden layers of 256 neurons, which maps the observation space $O_{t}$ to the action space $A_{t}$, and the neural network structure of the critical network is the same.
We generated a gait library that contains 125 distinct velocity commands. Each gait includes 26 different DoF positions or speeds, and the full-time step length for each periodic walking is 42. In Figure7, It can be observed that the closest gait space $G_{s}\subseteq\mathbb{R}^{n_{env}\times 42\times 26}$ is found for each environment based on its $x$ direction and $y$ direction velocity commands via our original gait library $G_{0}\subseteq\mathbb{R}^{125\times 42\times 26}$. At the same time, the original observation space $O_{t}\subseteq\mathbb{R}^{n_{env}\times 235}$ takes the corresponding dimensions to form a new control observation space $\hat{O}_{t}\subseteq\mathbb{R}^{n_{env}\times 26}$, so that the Mean Square of error(MSE) can be written as:
$L_{MSE}(O_{t})=E\Bigl{(}\bigl{\|}\hat{O}_{t}-\hat{G}_{t}\bigr{\|}\mid\hat{G}_{t}:=\{g_{i}\in G_{s}\mid argmin(\bigl{\|}g_{i}-o_{i}\bigr{\|})\}\Bigr{)}$ (3)
@@
      Compute advantage estimates $A_{t}$ and returns $R_{t}$ using GAE
@@ 4 Experimental Results and Analysis
In order to verify the performance ability of this locomotion control policy under various speed commands, we set speed tracking tests between $0.0\ m/s$ to $3.0\ m/s$ and compared it with the combination of Model Predictive Control (MPC) and Whole Body Controller (WBC)[9]. We define ”successful balance” as the ability to track the desired velocity continuously for 10 seconds without triggering any termination conditions. It was found that the balancing success rate at each speed, and the actual speed MSE at each speed are shown in the Table 1.
Command Velocity 0.2 0.4 0.6 1.0 1.5 2.2 3.0
@@
Ten trials are performed to obtain the experiment results. In each trial, 100 robots are running in parallel with the same controller in the simulation. From here, we can see that although MPC is a little better at slower speeds like $0.2\ m/s$, it will not be able to maintain balance at increased speeds. Also, the performance of both methods at $0.4\ m/s$, $0.6\ m/s$ and $1.2\ m/s$ $3.0\ m/s$ is shown in Fig.9
@@
The results indicate that our model has robust performance across a broad range of command velocities, given that the consistent MSE is observed for moderate-to-high speeds in Table.1. The MSE remains nearly constant as the command velocity increases, which means the control policy is well-matched with the gait library for velocities. At very low command velocities, such as $0.4m/s$, we observe a slightly worse velocity tracking performance compared to MPC; this issue presents an opportunity for further refinement. Moreover, when the command velocity reaches $3.0m/s$, the robot is unable to reach the desired speed, indicating an upper limit in the current gait formulation for high-speed locomotion.

### v2

@@ Abstract
Knee-less bipedal robots like SLIDER have the advantage of ultra-lightweight legs and improved walking energy efficiency compared to traditional humanoid robots. In this paper, we firstly introduce an improved hardware design of the SLIDER bipedal robot with new line-feet and more optimized mass distribution that enables higher locomotion speeds. Secondly, we propose an extended Hybrid Zero Dynamics (eHZD) method, which can be applied to prismatic joint robots like SLIDER. The eHZD method is then used to generate a library of gaits with varying reference velocities in an offline way. Thirdly, a Guided Deep Reinforcement Learning (DRL) algorithm is proposed to use the pre-generated library to create walking control policies in real-time. This approach allows us to combine the advantages of both HZD (for generating stable gaits with a full-dynamics model) and DRL (for real-time adaptive gait generation). The experimental results show that this approach achieves 150% higher walking velocity than the previous MPC-based approach.
Video: https://www.imperial.ac.uk/robot-intelligence/robots/slider/
Keywords: Bipedal Robot, Hybrid Zero Dynamics, Reinforcement Learning.
@@ 1 Introduction
Bipedal robots have promising applications in human-centered environments including assistive care and agile locomotion. The vast majority of bipedal robots, e.g., [1], are built with full revolute joints—a design approach that introduces singularity issues when the knees are fully extended during walking. Additionally, knee actuators contribute significant weight and increase the inertia of the legs. To address this challenge, knee-less prismatic-jointed robots such as SLIDER[2][3] have been optimized with lightweight leg designs, which in turn significantly reduce the mass and inertia during walking. Due to its inherent linear structure that functions as an impact cushion, [4] employed Bayesian optimization to demonstrate the energy savings of the prismatic joint across a range of frequencies.
However, similar studies have lacked system stability considerations over the entire dynamic system, i.e., the planar contact constraints of the entire supporting foot are needed to allow the robot to track reference trajectories during its swing phase[5]. A static, precisely matched reference trajectory does not imply that the trajectory can still be optimal in the dynamic system. In fact, the study of the long-horizon of the state of a robotic system shows that even if a reference trajectory is initially optimal under static conditions, its performance degrades over time due to the accumulation of dynamic effects, uncertainties, and disturbances[6]. It explain that why the long-horizon hybrid methods of learning and model-based control in manipulation[7] and the long-horizon walking trajectory optimisation such as Model Predictive Control (MPC)[5][8] have become increasingly popular. For robot locomotion, it is insufficient to design a controller solely on aligning the current state with the system’s linearized asymptotic stability point, especially when many robotic systems do not exhibit such a stability point at all[9].
@@ 2.1 Mass Reduction in Base Link and Child Joints
The subsequent child joints, including the yaw and roll joints, were redesigned with optimized, rigid 3D-printed parts featuring reinforcement ribs for enhanced strength in a compact form factor. Additionally, to minimize the radial load on the motor output shaft, both the roll and yaw axes adopt a coaxial mounting design. By integrating an additional bearing at the load’s distal end, the moment load on the motor’s internal output shaft bearing is reduced, improving motor lifespan and structural robustness. The overall assembly is illustrated in Fig. 2(b), with the name of each joint axis labeled.
@@ 2.2 Differential Capstan Drive for Foot and Prismatic Sliding Joints Actuation
To adapt SLIDER to the Capstan drive mechanism, a differential drive scheme was proposed, utilizing two identical actuators to control both the foot joint and the sliding joint simultaneously. An illustration of how this differential mechanism works is shown in the Fig. 3 below.
@@
As depicted, each leg assembly features a tightened steel cable on both sides symmetrically, which wraps around a pulley mounted on an actuator, with the endpoints bolted to the feet(shown in Fig. 4). Although only one foot is necessary, the cable must extend to the top and be anchored to an output lever to ensure proper functionality of the differential capstan drive throughout the entire sliding span. For simplicity, the lower foot design is mirrored at the top. Additionally, the linear motion constrain of the leg is handled by a LML9B metal linear rail, which is tough, smooth, and small enough for this application.
@@
With the reduced Degree of Freedom (DoF) in the foot joint and updated controller framework, only linear contact is required. Therefore the new foot design is much more simplified, which can be done using a 12mm aluminum extrusion. The new model weighs only 73.7g, which is just 80.9% of the previous version’s weight. Given the significant improvements outlined above, the new SLIDER design sets a benchmark for future hardware iterations and serves as a potential testing platform for the new control software framework. A complete CAD rendering and an assembled photo are shown in Figure 1 in the Introduction.
@@ 3.1 Extended Hybrid Zeros Dynamics
Here, the limit cycle is a closed trajectory in phase portrait. HZD cannot be directly applied to robots with prismatic joints, as the existence of a limit cycle is not guaranteed in such systems. Therefore, we propose the extended HZD method as shown in Fig. 5:
@@
Inside this figure, $S$ is the state space where the impact happened, and $Z$ is the invariant set based on Zero Dynamics assumption. $\Delta(\mathcal{S}\cap\mathcal{Z})$ means the discrete transformation when impact occurs. The left figure expresses the HZD applied in robot model with a virtual knee, and the right figure is the robot model with prismatic joints. The methods of mapping from left to right can be expressed by the following Fig. 6.
@@
In which the $q_{T}=[q_{\alpha}\ q_{\beta}\ q_{l}]^{T}$ express the position of hip pitch, ankle pitch, and sliding joint. $q^{\prime}_{T}=[q_{\alpha}\ q_{\beta}\ q_{l}]^{T}$, $q^{\prime}_{T}=[q^{\prime}_{\alpha}\ q^{\prime}_{\beta}\ q^{\prime}_{\theta}]^{T}$ express the position of hip pitch, ankle pitch and knee joint in virtual robot model. after the prismatic joint position transform $q_{l}=\sqrt{l_{1}^{2}+l_{2}^{2}+2l_{1}l_{2}cos(\theta)}$, equation. 1 can be applied to transform all state portrait from virtual knee model to the prismatic model.
Thus, once we have obtained the virtual limit cycle with the knee, we can obtain the approximation of the limit cycle in the state portrait of the ideal locomotion gait of the corresponding prismatic joint model. To implement HZD as described in [9], we set up the virtual constraints, adopting the configuration style of[10] to ensure the output matrix is full-rank.
$L(\theta)=\sum_{s\in S}\frac{1}{T}\,\Bigl{(}\sum_{t=1}^{T}\Bigl{\|}\tau_{s}\bigl{(}{q}_{t},{\dot{q}}_{t}\bigr{)}\Bigr{\|}^{2}\Bigr{)}$ (2)
We have used only the equation. 2 as the cost function to optimize the energy consumption. Here $T$ is the time period for each step, and $\tau$ is the torque output by the control law generated by the HZD method.
@@ 3.2 Guided DRL structure
In the simulation, We directly treat the upper foot and the lower foot as independent actuated joints while making the torque output of the upper foot always 0. We use the proximal policy optimization (PPO) algorithm with eHZD gait library for control policy training, and the whole training framework of Guided DRL is shown in Fig. 7:
@@
Here, the number of environments is $n_{env}$. The observation space is $O_{t}\subseteq\mathbb{R}^{n_{env}\times 235}$, which consists of CoM linear velocity space $v_{lin}\subseteq\mathbb{R}^{n_{env}\times 6}$, CoM angular velocity space $v_{ang}\subseteq\mathbb{R}^{n_{env}\times 6}$, weight projection space $g_{pro}\subseteq\mathbb{R}^{n_{env}\times 3}$, CoM velocity command space $C\subseteq\mathbb{R}^{n_{env}\times 4}$, actuated state position space $q_{dof}\subseteq\mathbb{R}^{n_{env}\times 12}$, actuated state velocity space $\dot{q}_{dof}\subseteq\mathbb{R}^{n_{env}\times 12}$, action space $A_{t}\subseteq\mathbb{R}^{n_{env}\times 12}$, and a height point cloud $H_{t}\subseteq\mathbb{R}^{n_{env}\times 180}$. The actions space is the desired position space for all actuated joints, which will be mapped to the final motor torque space by a PD controller. Here, the Actor network is composed of 3 hidden layers of 256 neurons, which maps the observation space $O_{t}$ to the action space $A_{t}$, and the neural network structure of the Critic network is the same.
We generated a gait library that contains 125 distinct velocity commands. Each gait includes 26 different DoF positions or speeds, and the full-time step length for each periodic walking is 42. In Figure7, It can be observed that the closest gait space $G_{s}\subseteq\mathbb{R}^{n_{env}\times 42\times 26}$ is found for each environment based on its $x$ direction and $y$ direction velocity commands via our original gait library $G_{0}\subseteq\mathbb{R}^{125\times 42\times 26}$. At the same time, the original observation space $O_{t}\subseteq\mathbb{R}^{n_{env}\times 235}$ takes the corresponding dimensions to form a new control observation space $\hat{O}_{t}\subseteq\mathbb{R}^{n_{env}\times 26}$, so that the negative Mean Square of error(MSE) can be written as:
$L_{MSE}(O_{t})=-E\Bigl{(}\bigl{\|}\hat{O}_{t}-\hat{G}_{t}\bigr{\|}\mid\hat{G}_{t}:=\{g_{i}\in G_{s}\mid argmin(\bigl{\|}g_{i}-o_{i}\bigr{\|})\}\Bigr{)}$ (3)
@@
Here $\tau_{t}$ is the torque output of each actuator, $v_{ang}$ and $v_{ang\_des}$ is the actual and desired angular velocity of main body, $v_{lin}$ and $v_{lin\_des}$ are the actual and desired CoM linear velocity, $H_{t}$ is the height of main body and $A_{t}$ is the position command for each joint at time $t$.
@@
      PPO: Compute advantage estimates $A_{t}$ and returns $R_{t}$ using GAE
@@ 4 Experimental Results and Analysis
In order to verify the performance ability of this locomotion control policy under various speed commands, we set speed tracking tests between $0.0\ m/s$ to $3.0\ m/s$ and compared it with the combination of Model Predictive Control (MPC) and Whole Body Controller (WBC)[5]. We define ”successful balance” as the ability to track the desired velocity continuously for 10 seconds without triggering any termination conditions. It was found that the balancing success rate at each speed, and the actual speed MSE at each speed are shown in the Table. 1.
Command Velocity 0.2 0.4 0.6 1.0 1.2 2.2 3.0
@@
Ten trials are performed to obtain the experiment results. In each trial, 100 robots are running in parallel with the same controller in the simulation. From here, we can see that although MPC is a little better at slower speeds like $0.2\ m/s$, it will not be able to maintain balance at increased speeds. Also, the performance of both methods at $0.4\ m/s$, $0.6\ m/s$ and $1.2\ m/s$ $3.0\ m/s$ is shown in Fig. 9.
@@
The results indicate that our model has robust performance across a broad range of command velocities, given that the consistent MSE is observed for moderate and high speeds in Table. 1. The MSE remains nearly constant as the command velocity increases, which means the control policy is well-matched with the gait library for velocities. At low command velocities, such as $0.4m/s$, we observe a slightly worse tracking performance compared to MPC; this issue presents an opportunity for further refinement. Moreover, when the command velocity reaches $3.0m/s$, the robot is unable to reach the desired speed, indicating an upper limit in the current gait formulation for high-speed locomotion.